# Simulation of Fermionic circuits using Majorana Propagation

### v1

@@ IV Length Truncation
In order to study Majorana length truncation analytically, we will consider unstructured Fermionic circuits with gates generated by Majorana monomials with $\norm{\boldsymbol{b}_{j}}_{1}$ an even number smaller than or equal to some $k^{\star}\ll N$. Unstructured here means that the generators are randomly sampled from the pool of such operators, $\{M_{\boldsymbol{b}}:\norm{\boldsymbol{b}}_{1}\leqslant k^{\star},\norm{\boldsymbol{b}}_{1}\mod 2=0\}$. This class of circuits can be regarded as a mathematically tractable, oversimplified model of certain relevant circuits, such as iteratively constructed variational circuits [51], which should suffice to elucidate the basic phenomena at play in Majorana Propagation for typical Fermionic circuits.
@@
The first important point motivating Majorana length truncation is the fact that most high-length terms do not contribute to the expectation value $\langle H\rangle$. As discussed above, only paired Majorana monomials yield non-zero values when traced with a Fock basis state. The probability $P(w)$ for a randomly chosen $w$-monomial to be paired, is given by
@@
and as depicted in Fig. 2, is very small unless $w$ is close to 00 or to $2N$. This is stated more precisely in the following theorem (see Appendix C.1 for proof).
Theorem 1 (The contribution of high length Majorana monomials are exponentially suppressed).
@@
Theorem 2 (Suppression of backflow).
Consider the propagation of a length $w$ Majorana monomial $M$ under a length $k$ generator of rotation $M^{\prime}$ as per Eq. (5). We assume that $[M^{\prime},M]\neq 0$ such that the propagation causes branching and we denote the length of the new Majorana monomial, $M^{\prime}M$, as $w^{\prime}$. Let $P_{+}$ and $P_{-}$ denote the probabilities that the propagation increases ($w^{\prime}>w$) and decreases ($w<w^{\prime}$) operator length. The following statements hold:
@@
The algorithm introduced in this work takes advantage of this dynamics. As explained above, the monomials clustering around $w=N$ contribute exponentially little to the trace with the Fock basis state, so they can be neglected. Due to the inversion in the branching imbalance for $w>N$, the terms with $w\approx 2N$ are never populated, so they need not be considered in the simulation. Moreover, since backflow probability throughout the dynamics is very small, there is no need to account for high-length monomials until they reach the $w\approx N$ regime. Hence, they can be neglected early on, as soon as they leave the low-length subspace.
V Implementation
Figure 3: Majorana Propagation implementation for ground states. Top: The blue curve depicts the error in the estimated energy as a function of monomial length cut-off $w^{\star}$ for a 28-mode MP-simulated ADAPT-VQE circuit containing 22 Fermionic gates, computed by comparison against statevector simulation (SV). The error of the MP simulation decreases exponentially from below chemical precision (1.6 mHa) for $w^{\star}=4$ to around numerical precision for $w^{\star}=10$. The green curve shows the same quantity for 5 circuits obtained by uniformly randomizing angles in the range $[-0.1,0.1]$. The shaded area indicates the maximal and minimal errors attained. Bottom: The ADAPT-VQE inspired algorithm is run classically using MP to simulate the circuits for increasing problem sizes (solid lines). The $x$-axis reflects the number of Fermionic gates in the circuit. The $y$-axis value is the error in the energy, computed with MP with $w^{\star}=6$, divided by the spectral gap of the Hamiltonian, which bounds the lack of overlap with the ground state. For 28 modes we verify the accuracy of the MP simulation by computing the equivalent error via a SV simulation (yellow markers). SV simulation for the larger problems was not tractable. Only every 5th marker is included to improve readability.
Number Majorana Propagation Qiskit CUDA-Q CUDA-Q
of modes pre-pro. estim. SV SV MPS
28 0.32 s 0.0023 s 3.71 h 12.18 min 32.15 min
40 7.86 s 0.0461 s 10.10 h
52 104.93 s 0.2145 s $>$ 24 h
In Fig. 3 (top), we plot the error in the energy estimated with MP as a function of the length cut-off $w^{\star}$ for an ADAPT-VQE circuit comprising 22 Fermionic gates—equivalent to a 616-CNOT circuit on all-to-all connectivity (see Appendix D)—optimized for the 28-mode active space Hamiltonian. The error $\varepsilon$ is computed as $\varepsilon=|E_{\rm MP}(w^{\star})-E_{\rm sv}|$, where $E_{\rm sv}$ is the exact energy (up to numerical precision) obtained via statevector simulation. Remarkably, even for $w^{\star}=4$, MP produces an estimate with an error below chemical precision (1.6 mHa), with the error decreasing approximately exponentially as a function of the cut-off to nearly numerical precision for $w^{\star}=10$. A similar behaviour is observed when the angles are randomized within the parameter range in the optimized circuit, although the decay in $w^{\star}$ is slower.
In many applications, the relevant figure of merit quantifying the quality of a reference state $|\phi\rangle$ is its overlap with the ground state $|\langle\phi|\psi_{0}\rangle|$. Given the error in the energy, $|E-E_{0}|$, the overlap can be lower-bounded if the spectral gap of the Hamiltonian is known (see Appendix E). Thus, using the spectral gap computed with DMRG, we can upper-bound the state error, which we define as $1-|\langle\phi|\psi_{0}\rangle|$. The bottom plot in Fig. 3 shows the upper bound to the state error as a function of the number of Fermionic gates in the circuit as the adaptive training progresses. All cases show a rapid convergence, reaching an overlap of at least 97% with the ground state with 22, 102, and 300 Fermionic gates for the three active space sizes (the simulations were manually interrupted upon reaching a absolute error of 10 mHa). When mapped to qubit space and transpiled for an all-to-all connectivity quantum computer, the equivalent circuits contain 616, 3470, and 13,676 CNOTs, respectively, which highlights the complexity of the obtained states.
Finally, Table 1 reports run-time benchmarks for energy estimation with MP and other methods: two statevector simulators (Qiskit and CUDA-Q), as well as CUDA-Q MPS simulator, a tensor network-based circuit simulator. The circuits used are the converged ADAPT-VQE circuits from Fig. 3 (bottom), transpiled to qubit space, resulting in the CNOT counts reported above. While statevector simulators are exact (up to numerical precision), they scale exponentially. Thus, only the 28-mode circuit could be simulated with these simulators with reasonable resources. MP and CUDA-Q MPS, on the other hand, produce approximate estimates but run in polynomial time. Details on the implementation of these methods can be found in Appendix D. Interestingly, in addition to simulating these circuits considerably faster than CUDA-Q MPS, the 28-mode result is also nearly two orders of magnitude more accurate, approximating the statevector energy within $\approx 1.39\cdot 10^{-2}$ mHa, as opposed to the $\approx 2.71$ mHa error in the MPS simulation.
VI Discussion
Our analytical results portray MP as a very natural framework for the simulation of Fermionic circuits. While originally inspired by PP, MP is not equivalent to the former. MP leverages newly identified phenomena in the behaviour of basis monomials, which are not related to Pauli weight. The underlying assumption in our analysis, namely that circuits are unstructured, with gate generators randomly chosen from a set of bounded-length monomials, seems particularly appropriate to model iteratively constructed ground state approximating circuits. Our analysis also sheds light on the working principles behind approaches like Majorana-length-damping in tensor network methods for Fermionic systems [58].
Our numerical results substantiate that MP can be used to classically simulate ADAPT-VQE to produce circuits with significant overlap with the ground state in strongly correlated systems. In our simulations, MP was able to outperform state-of-the-art, tensor network circuit simulators. It would be interesting to compare its performance to more recently proposed approximate simulation methods for Fermionic circuits, like Refs. [59, 60, 61, 62]. Since MP does not rely on the sparsity of the wave-function in the Fock basis, and since single-excitation gates do not require additional truncation in MP, MP may produce more accurate estimations for circuits with many such excitations leading to dense states. Similarly, we stress that MP is strictly more powerful and efficient than ‘gsim’ approaches [63, 64, 65, 66, 67] as the basis of the algebra is effectively generated ‘on the fly’ during the simulation (rather than in advance) and so is not restricted to free Fermionic models with polynomially scaling Lie algebras.
@@ Acknowledgements
GGP conceived the method. AM implemented the method. ZH and GGP formalized the theory and wrote the first version of the manuscript. AM, OS, AN, RC, and AG implemented and executed the ADAPT-VQE simulations under AG’s direction. AM, OS, and AG produced the figures. All authors participated to discussions and to the writing of the manuscript.
@@ A.1 Introduction to fermionic systems
Mathematically, the $N$-mode Fermionic system is equivalent to the $N$-dimensional Fock space $\mathcal{F}(\mathbb{C}^{N})$, a $2^{N}$-dimensional Hilbert space spanned by the so-called Fock basis. The operators defined above allow us to define the basis as follows. First, the Fermionic vacuum |$\ket{\text{vac}_{\text{f}}}$ is defined to be the unique vector such that $a_{j}\ket{\text{vac}_{\text{f}}}=0$ for all $j=1,\dots,N$. The remaining basis elements, namely Fock states, can be constructed by considering all possible combinations of occupation numbers $n_{j}\in\{0,1\}$:
|$\ket{n_{1}n_{2}\dots n_{N}}\coloneqq\prod_{j=1}^{N}(a_{j}^{\dagger})^{n_{j}}\ket{\text{vac}_{\text{f}}}.$ (15)
@@
Thus if $w$ or $w-1$ is a multiple of 4 then $m_{x_{1}}...\,m_{x_{w}}$ is already Hermitian and $r=0$, but otherwise $m_{x_{1}}...\,m_{x_{w}}$ is antihermitian and $r=1$ so that the factor of $\mathrm{i}$ makes the monomial Hermitian.
@@ A.2 Algorithm overview
where $\varrho=\outerproduct{n_{1}n_{2}\dots n_{N}}{n_{1}n_{2}\dots n_{N}}$ |, with $n_{i}\in\{0,1\}$, is a Fock basis state. The circuit
@@
Our Majorana Propagation method, analogously to Pauli propagation algorithms, works in the Heisenberg picture where the initial operator (here the observable $H$) is often sparse in the Majorana basis. In this basis, the operator is back-propagated through the circuit, i.e., we compute $U(\boldsymbol{\theta})^{\dagger}HU(\boldsymbol{\theta})$, and finally overlap with the initial state $\varrho$. It is well known that the effect of applying a unitary gate generated by a single Pauli string $P$ to another Pauli string $Q$ is given by,
@@ Appendix B Coefficient and small angle truncations
$t\in\mathcal{O}\left(N\cdot L^{\log{\left(\frac{1}{\epsilon^{2}\delta}\right)}}\right)\;,$ (29)
@@
$t\in\mathcal{O}\left(N\cdot L^{\log{\left(\frac{1}{\epsilon}\right)}}\right)\;$ ) (31)
@@
Appendix C Length Truncation
@@ C.1 The contribution of typical high length Majorana monomials are exponentially suppressed
[…]
@@
Thus if we set $w/2=cN$ ($10<c<$ 1) we have that
@@
where $p=ec$. Hence if $c<1/e$, corresponding to $w<2N/e<0.74N$, we have that $10<p<$ 1 and thus $P(w)=r^{-N}$ with $r=p^{-c}>1$. To quantify the scaling of $P(w)$ in the region $0.74N\leqslant w\leqslant N$ we note that $P(w)$ decreases monotonically with $w$ to its minimal value at $P(N)$ and hence
@@
with $r=p^{-c}$ for some $10<p<$ 1 for any $w=Nc$, $10<c<$ 1. Finally, since $P(w)$ is symmetric with respect to $w=N$, $P(w)=P(2N-w)$, the result applies to any $w=2Nc$, as claimed.
@@ C.2 The probability of backflow is suppressed
Consider the propagation of a length $w$ Majorana monomial $M$ under a length $k$ generator of rotation $M^{\prime}$ as per Eq. (25). We assume that $[M^{\prime},M]\neq 0$ such that the propagation causes branching and we denote the length of the new Majorana monomial, $M^{\prime}M$, as $w^{\prime}$. Let $P_{+}$ and $P_{-}$ denote the probabilities that the propagation increases ($w^{\prime}>w$) and decreases ($w<w^{\prime}$) operator length. The following statements hold:
@@
Part 1) Let us denote the number of overlapping Majorana operators between $M$ and $M^{\prime}$ as $s$. The commutator $[M,M^{\prime}]\neq 0$, i.e., the propagation only causes branching, if and only if the number of Majorana operators common to both monomials is odd. Thus in what follows we can assume $s$ is odd.
@@
    Choose the remaining $w-s$ non-overlapping elements: The remaining $w-s$ elements of the second string must come from the $2N-k$ elements that are not in the fixed string. This selection can be made in $\binom{2N-k}{w-s}$ ways.
@@
We can assume that $s<k$ (if $s=k$ then the ratio in Eq. (58) evaluates to 0 and so is less than 1 as required) and $s<w$ (as we already assumed $k\leqslant w$). We now further assume that $02N-k-w+s+1>$ 0 for all $s$, which holds if

### v3

@@ Discussion.
In order to study Majorana length truncation analytically, we consider unstructured Fermionic circuits with gates generated by Majorana monomials with $\norm{\boldsymbol{b}_{j}}_{1}$ an even number smaller than or equal to some $k^{\star}\ll N$. Unstructured here means that the generators are randomly sampled from the pool of such operators, $\{M_{\boldsymbol{b}}:\norm{\boldsymbol{b}}_{1}\leqslant k^{\star},\norm{\boldsymbol{b}}_{1}\mod 2=0\}$.
@@
The first important point motivating Majorana length truncation is the fact that most high-length terms do not contribute to the expectation value $\langle H\rangle$. As discussed previously, only paired Majorana monomials yield non-zero expectation values when traced with a Fock basis state. The probability $P(w)$ for a randomly chosen $w$-monomial to be paired, is given by
@@
and as depicted in Fig. 4, is very small unless $w$ is close to 0 or to $2N$. This is stated more precisely in the following theorem (see Appendix C.1 for proof).
Proposition 3 (The contribution of high length Majorana monomials are exponentially suppressed).
@@
Proposition 4 (Suppression of backflow).
Consider the propagation of a length $w$ Majorana monomial $M$ under a length $k$ generator of rotation $M^{\prime}$ as per Eq. (5). We assume that $M^{\prime},M]\neq 0$ such that the propagation causes branching and we denote the length of the new Majorana monomial, $M^{\prime}M$, as $w^{\prime}$. Let $p_{+}$ and $p_{-}$ denote the probabilities that the propagation increases ($w^{\prime}>w$) and decreases ($w<w^{\prime}$) operator length. The following statements hold:
@@
The algorithm introduced in this work takes advantage of this dynamics. As explained above, the monomials clustering around $w=N$ contribute exponentially little to the trace with the Fock basis state, so they can be neglected. Due to the inversion in the branching imbalance for $w>N$, the terms with $w\approx 2N$ are never populated, so they need not be considered in the simulation. Moreover, since backflow probability throughout the dynamics is very small, they can be neglected early on, as soon as they leave the low-length subspace.
We finally need to account for the fact that while backflow is suppressed in theory the contribution of many unlikely backflows could add up to a substantial contribution to the expectation value. However, in practice this does not occur as the total two norm of the evolved operator is conserved under unitary evolutions and so the coefficients of the terms that do backflow are small.
In Appendix D we tie together these intuitions to prove that the average error of a length-truncated Majorana propagation simulation is exponentially suppressed in truncation length $w_{0}$. To do so we employ a Markov chain analysis that is conceptually rather different to the proof strategies employed to analyze path propagation methods previously [10]. In particular it allows us to first obtain an efficient classical algorithm for computing the mean squared error exactly.
Theorem 5.
The mean squared error $\operatorname{\mathbb{E}}((f_{L}-f_{L}^{(w_{0})})^{2})$ for a random Majorana circuit as defined in Def. 4 can be computed exactly classically in polynomial time.
It is this algorithm that we use to plot the exact average error in Fig. 2. We then proceed to bound the error. Crucially, in contrast to proofs for PP [10], we obtain a depth dependent error bound as shown in Theorem 1 and plotted in Fig. 2. In fact, our algorithm can also be applied more generally for computing the average of the expectation value. It would be interesting to investigate whether this strategy could be applied to propagation methods in other contexts to compute coarse-grained properties and/or derive tighter bounds.
@@ Acknowledgements
GGP conceived the method. AM implemented the method. ZH and GGP formalized the initial theory and wrote the first version of the manuscript. JF conducted the end-to-end error analysis under the supervision of ZH and ZZ. AM, OS, AN, RC, and AG implemented and executed the ADAPT-VQE simulations under AG’s direction. AM, OS, and AG produced the figures. All authors participated to discussions and to the writing of the manuscript.
@@ A.1 Introduction to fermionic systems
[…]
@@
Thus if $w$ or w1w-1 is a multiple of 4 then $m_{x_{1}}...\,m_{x_{w}}$ is already Hermitian and $r=0$, but otherwise $m_{x_{1}}...\,m_{x_{w}}$ is antihermitian and $r=1$ so that the factor of $\mathrm{i}$ makes the monomial Hermitian.
@@ A.2 Algorithm overview
where $|\varrho=\outerproduct{n_{1}n_{2}\dots n_{N}}{n_{1}n_{2}\dots n_{N}}$, with $n_{i}\in\{0,1\}$, is a Fock basis state. The circuit
@@
Our Majorana Propagation method, analogously to Pauli propagation algorithms, works in the Heisenberg picture where the initial operator (here the observable $H$) is often sparse in the Majorana basis. In this basis, the operator is back-propagated through the circuit, i.e., we compute $U(\boldsymbol{\theta})^{\dagger}HU(\boldsymbol{\theta})$, and finally overlap with the initial state $\varrho$. It is well known that the effect of applying a unitary gate generated by a single Pauli string PP to another Pauli string $Q$ is given by,
@@ Appendix B Coefficient and small angle truncations
[…]
@@
[…]
@@
Appendix C Length Truncation: Intuition
@@ C.1 The contribution of typical high length Majorana monomials are exponentially suppressed
In this subsection we substantiate the claim that the higher length Majorana monomials $M_{\boldsymbol{b}}$, that arise during a propagation, are more likely to annihilate when we compute their overlap with the initial Fock state. Intuitively, this can be seen by supposing that, due to the circuit’s complexity, all high-length $w$-monomials carry a similar length. There are $\binom{2N}{w}$ such monomials. The number of these that are paired is $\binom{N}{w/2}$. The probability $P(w)$ that any randomly chosen length $w$ Monomial $M_{\boldsymbol{b}}$ is paired (and therefore Tr[$\Tr[\varrho M_{\boldsymbol{b}}]\neq 0$) is given by
@@
Thus if we set $w/2=cN$ ($0<c<1$) we have that
@@
where $p=ec$. Hence if $c<1/e$, corresponding to $w<2N/e<0.74N$, we have that $0<p<1$ and thus $P(w)=r^{-N}$ with $r=p^{-c}>1$. To quantify the scaling of $P(w)$ in the region $0.74N\leqslant w\leqslant N$ we note that $P(w)$ decreases monotonically with $w$ to its minimal value at $P(N)$ and hence
@@
with $r=p^{-c}$ for some $0<p<1$ for any $w=Nc$, $0<c<1$. Finally, since $P(w)$ is symmetric with respect to $w=N$, P(w)=P(2Nw)P(w)=P(2N-w), the result applies to any $w=2Nc$, as claimed.
@@ C.2 The probability of backflow is suppressed
Consider the propagation of a length $w$ Majorana monomial $M$ under a length $k$ generator of rotation $M^{\prime}$ as per Eq. (29). We assume that $M^{\prime},M]\neq 0$ such that the propagation causes branching and we denote the length of the new Majorana monomial, $M^{\prime}M$, as $w^{\prime}$. Let $p_{+}$ and $p_{-}$ denote the probabilities that the propagation increases ($w^{\prime}>w$) and decreases ($w<w^{\prime}$) operator length. The following statements hold:
@@
Part 1) Let us denote the number of overlapping Majorana operators between $M$ and $M^{\prime}$ as $s$. The commutator $M,M^{\prime}]\neq 0$, i.e., the propagation only causes branching, if and only if the number of Majorana operators common to both monomials is odd. Thus in what follows we can assume $s$ is odd.
@@
    Choose the remaining wsw-s non-overlapping elements: The remaining wsw-s elements of the second string must come from the 2Nk2N-k elements that are not in the fixed string. This selection can be made in $\binom{2N-k}{w-s}$ ways.
@@
We can assume that $s<k$ (if $s=k$ then the ratio in Eq. (62) evaluates to 0 and so is less than 1 as required) and $s<w$ (as we already assumed $k\leqslant w$). We now further assume that 2Nkw+s+1>02N-k-w+s+1>0 for all $s$, which holds if